10th term of the progression -4-1+2+5+------------

Grade:

2 Answers

Godfrey Classic Prince
11 years ago

Dear jayku kumar,

The Formula for finding the nth term in Arithmetic Progression is given by

an = a1 + (n-1)d

where,

an is the value of the nth term

a1 is the value of the first term

n is the nth term

d is the difference 

 

In this case, a1 = -4, d = (a2 -a1) = 3, n=10

Therefore substituting these values,we get the value of the nth term(an)

Therefore, an = -4 + (10-1)3

                   = -4 +(9)3

                   = -4+ 27

                   = 23

Hence the 10th term is 23.

 

Hope this helped you immensely...!!

All the Very Best & Good Luck to you ...!!

 

Best Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-M

Please approve my answer if you liked it by clicking on "Yes" given below..!!Smile


 

RAJORSHI PAUL
11 years ago

the given sequence is an AP with a1= -4 and d= 3. We have find out the 10th term of this sequence.

So an = a1 + (n-1)d. Here n=10. 

So a10= -4 + 9×3 =23

Hence the 10th term of this sequence is 23

 

RAJORSHI PAUL

Dept of Electrical Engg

IIT Kharagpur